Tech Staff Meeting Notes November 13, 2008

Current Top Priorities are: Station maintenance

The Christmas party is 11-1 on December 4. Michelle Smith is organizing it.

We need to make sure we do a site visit form each time we visit a site.

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

The task list updated February 6, 2008. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.



Jim's Priorities

Several of the stations can get improved telemetry by transmitting directly to the node. The original configuration changed enough so that there is room in the band. Jim will investigate.

Holly's list of problem stations:

ISIS Oxidation Problem


Muted amplitude

(PLAR note from last time retained only as a reminder for Mitch) The problem with PLAR appears to be cultural. Jim surmises there is a large motor or something nearby. The data are still useable but the noise threshold is about 4 times higher than normal and very peaked at 5 and 6 Hz.

Notes from John on the ET trip 11/3-7:
Monday: WSNC - Completely installed in one day!
Tuesday: RBNC - Added 2nd solar panel, repaired unknown problem with Z VCO, Installed WSNC receiver and mux card. Balanced signals. NC Node - Moved rack to new building
Wednesday: NC Node - Spliced cables at old building and West receiver box. Installed 2 new cables from West receiver box to new rack location. Grounded everything as it enters building. Installed new GPS pod and GPS receiver. Installed new fiber from rack to ethernet connection in the radio station's transmitter room. Tested rack and noted problem stations.
Thursday: CCNC - Arrived expecting to find overdriven transmitter, but found that the pit box was on its side with a significant amount of water in it. Repaired pit box, hardened it with mortar mix, devised new entry method for balanced line, replaced old baluns with new style BABs. Transmitter itself had somehow become misaligned and Jim was able to re-adjust the deviation to the proper levels without the aid of the proper instruments. TZTN - Swapped out gateway and got station back up.
Friday: fgtn - Arrived to find problem with LRVA was not there. LRVA - Found that new baluns have a larger than expected voltage drop across the balanced line, Replaced the new BABs with the old ones, replaced battery and replaced misaligned Z VCO card.

CCRT - Needs a rebuild after a bad lightning strike.

GFM - Fixed the bad channel but still needs some work to repair storm damage. The NC state park service is purchasing GFM. We already have a couple permissions from them and John initiated the paperwork for the GFM permission.

GMG - Need battery box solar reg mounting plate and need to fuse the battery.


Second Tier Priorities

Replace batteries in all XLAK clocks. Need to be careful so that earthworm doesn't see time jump around. (See Jim's priorities)

Most of the strongmotion stations can run at 100 sps without any modifications. We will install a router on the few that have problems (RDST will be the test case). We will also migrate the acquisition software from gcf2ew on anssworm to scream/scream2ew on a linux or ms server. The latter will facilitate stations maintenance without interrupting the data stream. (See Jim's priorities).

All CMG5TD sensors need to have the latest firmware (59914) installed and trigger download verified. Also need to adjust the trigger level as warranted by the local noise. The nominal level is 7000 "micro-g" which is really counts and translates to about 1% g (nominal conversion is 130,000 count/m/s/s). Note that some DAS's are fussy and you may need to set them to filing mode and then to dual to convince them to save triggers. Note that we're triggering on the 10sps channel which needs a fairly long period signal to trigger. A tap on the sensor may not be enough. Instead the bolts may need to be loosened to move the sensor and cause a trigger. All the west-side stations are done. We'll do the east side ones when everything is in scream and when John is able to make a site visit if something hangs (he's scheduled to be in Charleston early december).

Need a new bottom air conditioner for mkta. Need new RF preamps, radio align, etc, etc at MKTA.
Need to follow up the direct assault on the embryophyte hordes with chemical warfare.
Need to swap out the old node computers at AZTN. Will XP do netbui with the bridge recorders (if so we could lose the windows 98 box at aztn)?

MACC went deaf and dumb. James is trying to see if we can enlist Greg's help in retrieving/installing hardware.
Clearly, some of these items are lower priority.
CVMS. The temp cover James made for the GPS enclosure is gone. We've been talking about making a combined cover for the CVMS and CVTN enclosures. he two are side-by-side and could easily be covered by something like a lean-to.
PTGV. The temp enclosure cover is still in place. Chris has been talking about making a more permanent one similar to the louvered cover we have at CERI and at STLE. Chris and Bob are still working out the details, namely how much it might cost. Chris also mentioned replacing the current enclosure with a larger one. We have a larger enclosure available in the Expansion Space should this come about.
CJTR. We're still waiting for a definite "yes" or "no" from Haydar about putting a PC there to call CJTR locally. Steve's going to get me an old GX1 to make into a Linux box for the job.
CSMS. The previous comments James made about a new department taking of the CSMS facilities, where our station is, are false. They're still waiting for money to start construction on the new CSMS buildings. We're stuck with them for a while longer. According to the locals, wi-fi is prohibited on-base (even though it's apparently available elsewhere on the base). There is no wi-fi close enough for us to use.
All H-beam monument Sites (CVMS, HCES, STLE, MAIR, MCTY, NWCC, RLAP, PTGV, PIGT). All sites using a deep-driven pile need to have some cover placed over the PVC surrounding the monument at the service. Chris and James discussed some options and he's looking at prices. Priority goes to schools or places where small kids could get twisted ankles or other injuries (HCES, CVMS, RLAP).
All Sites except NMKM, LCHS, CJTR, and MACC. All sites may need their enclosures re-painted, preferably white. MCTY, CVMS, HCES, and NWCC currently have yellow enclosures. Painting them white might help them stay cool.
MCTY. Struck by lightning. Receiver is at Ashtech for repair. Ashtech wants the antenna and choke ring too.


Fort Payne - John found a good spot with decent odds of a permission at Desoto state park in Alabama. It will have to repeat through a Georgia fire tower and John is investigating the permit there.

The map shows stations with yellow triangles being new ones.
WSSR8done (now WSNC)
Maryville8moved to northeast AL
SMNC4done (see note above)
BHT8move (eventually)

The recorder disks should be reformatted as a routine maintenance task next time access is available. The constant read/write to the fat file systems lead to fragmentation. Don't ya just love dos?
Lost contact with recorders 1, 3, and 6. Number 5 is unstable. Can reset all 7 recorders which implies the equipment enclosure at each node has power and that at least the reset radio and telemetry are working.

Jim and Steve participated in the TDOT/TransCore meeting regarding the Boswell work. It now looks like we won't be on their wireless and won't need the T1 channel bank hardware. More likely we'll need to telemeter to a point at 2nd street and get on the fiber directly. If that's true it might be routable addresses in which case we can connect directly via CERI without anything at Boswell or additions to Richardson. They are committed to sharing the network but are going full bore at Boswell and prefer to not be distracted. We're in a wait and see until that's done possibly around October 6.

Mud Island vertical array is operational. Not really, functional is a better word until we can get it online to qc/vet data. Trigger levels on the strongmo need to be adjusted because the flash fills in about two weeks. One dm24 is at Guralp for repair. The sensor is in 3904/215E.

CVTN optical isolation. Need to upgrade to year 3 hardware. Hardware is ready.

Steve needs to speak with Dean Kemp at NWS. They are willing to let us piggyback on their cable internet (not

Lenox Agreement. Jim obtained contact info for the person Mitch needs to call>

FDL. This is the FDL Policy. .
Horton and Langston deployed the CMG6TD's to record Gustav and found that 3 of them were not operational. Steve and Dave will get together to go over all of them to "check-in" the ones that are ok and to assess what the others need to make them operational.
Dave fixed the bad cables but still needs a way to waterproof the usb connectors.
Mitch will arrange some CMG6TD review with Horton and tech staff.


As Time Permits

There is SOH info superimposed on the pII gain channel. Jim and John will get the codes to Mitch who will investigate the feasibility of decoding software.

PEBM and HENM sometimes have telemetry fades. Should reevaluate comms.

Need to gather up a contact list (e.g. everyone's home phone, power companies, etc) and put it on webworm (password protected).

Route Pat's alarms into the earthworm digitizer. They are a flat -5/0/+5 voltage that monitors a state change, it should then be easy to add them into the env_alarm channels.

CPCT - Re-work lighting and pest chaser. John has a pest chaser and will use some landscape or pool hardware for the lighting.

SWET - Need to seal the hinges on the aluminum Hoffman boxes with flashing tape to keep the water out. Done, need to check it. Is this the permanent solution or a temporary fix?

3876 Power The power for the new lighting on the east side of the general parking lot to the west of CERI is coming from 3876. Chris will check it out to assess the added lightning hazard and see if anything can be done about it. Might be hard to do without the wiring diagram. Chris will make one if needed as time permits.

3890 Power Chris is going to make a diagram for 3890. Good luck son.

Inventory Mitch needs to look at this and determine where we are and what needs to be done to make this operational.



Withers needs to determine current orientation for MPHB so we can pull it and orient correctly.

Pink Palace is working with Gary and Michelle on a new perc-like display area. We'll pull the receiver etc, when they start pulling out the old display (or sometime earlier if an opportunity arises).

Jim will switch the FCC licenses from Greg's name to Mitch. Done.

The Transportable Array (TA) component of the U.S. Array observatory of the Earthscope Facility will be installed in the NMSZ in 2011 and permissioning in AR and MO will be the summer of 2009. We wish to provide a list of CERI partner landowners to USArray that are willing to host TA stations and, if possible, a list of landowners who are not willing to host stations. Mitch will supply Chris with a list of target locations (approximately 70km grid).

Mitch briefly reviewed the desired 106x6 network of desireable TA stations to adopt. These are the 106 TA stations within the ANSS RSN footprint for the NSMZ and WVSZ and the ETSZ and CSZ. The "by 6" implies three components for broadband and three of strongmotion. Considerable funds must be secured for the capitalization costs and for the long term operation and maintenance.